# Signed Networks: theory, methods, and applications

### v1

@@ 5.2.2 Data-driven approaches
A third approach to gathering empirical signed networks involves extracting data from online platforms. This strategy has gained popularity thanks to the ease of data collection and the ability to access systems involving thousands of agents. One of the earliest studies in this direction is [kunegisSlashdotZooMining2009], which analyses the Slashdot Zoo network but does not focus on temporal evolution. A more complete dynamic analysis appears in [szellMeasuringSocialDynamics2010, szellMultirelationalOrganizationLargescale2010, szellSOCIALDYNAMICSLARGESCALE2012], which examines the multilayer network of Pardus, a massive multiplayer online game with over 300,000 nodes. The structural analysis, though static, already points to key mechanisms in the network. They determine three different ways in which positive and negative links differ: first, positive links are highly reciprocal, whereas negative links are not; second, negative links follow a power-law degree distribution, unlike positive ones; and finally, positive layers exhibit higher clustering. In addition, the hubs in the positive and negative subgraphs are uncorrelated. The network also appears more balanced than expected by chance. In their analysis of network dynamics, the authors examine the network evolution over three years. Once again, they find qualitatively different behaviour in the link dynamics depending on their sign: negative links follow preferential attachment, whereas positive ones do not. In addition, they distinguish between two types of negative links: private enemies, which are typically local and uncorrelated with balance, and public enemies, who act as negative hubs and are often jointly targeted by friends, thus playing a central role in the formation of balanced triads. A follow-up study [sadilekAsocialBalanceHow2018] develops a generative model to simulate Pardus’ dynamics. They show that, while the dynamics of positive links are relatively autonomous, the dynamics of negative links are tightly coupled to the positive layer. Other examples of such analyses are [liModelingOnlineSocial2018, askarisichaniStructuralBalanceEmerges2019, liuEvolutionStructuralBalance2020].
@@ 5.2.3 Mechanisms governing signed network dynamics
As we have seen, many empirical analyses agree that the network evolution is shaped by mechanisms beyond structural balance. Identifying these mechanisms has proven difficult, partly because they are often specific to the system under study and partly because they interact in complex, nontrivial ways. In addition, common simplifications such as removing link directionality can obscure important aspects of the dynamics. More fundamentally, the mechanisms we have covered so far, which we will name endogenous mechanisms, are just one part of the picture. In general, we can group the factors driving network change into three types [riveraDynamicsDyadsSocial2010]. (i) Endogenous mechanisms: some links appear or disappear by the simple presence of other links in the network, so that the current structure of the network constrains and modifies its future structure. These mechanisms include balance, reciprocity, or triadic closure. (ii) Node-based attributes that are not encoded in the network structure but affect the way the network evolves over time. Examples include personality, social status, or shared interests. (iii) Contextual and external factors, coming from the environment in which the network is embedded. These include institutional settings, cultural norms, or platform algorithms.
@@
    Homophily: the tendency to form positive links with others who share similar characteristics, such as interests, beliefs, or social status [riveraDynamicsDyadsSocial2010]. This mechanism is closely related to, and often hard to separate from, influence, which is the tendency for connected individuals to become more similar over time. Homophily can have three variations: homophobia/heterophily/heterophobia, which would be the tendency to form negative/positive/negative links with those nodes with similar/different/different characteristics, respectively. These variations can be generalised with the concept of assortativity.
@@
    Reciprocity: the tendency for directed links to become bidirectional. This results in an overrepresentation of mutual links compared to a null model. Moreover, when asymmetrical links do not become reciprocal, they tend to be unstable and eventually disappear [riveraDynamicsDyadsSocial2010].
@@
    Transitivity and clustering: the tendency to complete the connection $p\rightarrow q$ given the existing connections $p\to o$ and $o\to q$. In undirected networks, this mechanism is known as clustering or triadic closure, the tendency to close open triangles. This effect seems to be cumulative and grows with the number of common neighbours shared by two nodes. [riveraDynamicsDyadsSocial2010].
@@
    Differential popularity: the tendency of networks to display heterogeneous degree distributions for both positive and negative links [riveraDynamicsDyadsSocial2010]. This mechanism includes two distinct dimensions: activity, which refers to variation in the number of outgoing links, and popularity, which refers to variation in the number of incoming links. At a dynamic level, differential popularity can appear by means of preferential attachment or detachment, a mechanism through which the rates of gain or loss of links depend on the degree of the node [barabasiEmergenceScalingRandom1999a]. It should be noted that the different ideas encompassed within the generic term “differential popularity” may be related to each other, giving rise to theories such as visibility theory, popularity theory, activity theory or karma theory. All of these possibilities are collected in [yapWhyDoesEverybody2015].
@@
    Resource allocation: forming and maintaining links requires time, attention, and cognitive effort, which are all limited. This constraint shapes how many links a person can have and how these social bonds are structured [powellOrbitalPrefrontalCortex2012, tamaritCognitiveResourceAllocation2018, tamaritDunbarCirclesContinuous2022]. These resource constraints shape the degree distributions for both positive and negative links, and can explain why negative links are often unstable (conflict is costly). It also helps explain why seemingly stable structures can disappear: when links dissolve, the freed resources can be reinvested in new connections, and conversely, existing links may need to be dropped to make room for others.
@@
    Memory: while many models assume Markovian dynamics (i.e., memoryless evolution), empirical evidence suggests that past interactions can influence future ones. For instance, nodes may be more likely to re-establish past positive links or avoid reconnecting after a negative link has disappeared. A similar reasoning has been suggested to influence the process of link removal [kleinbaumReorganizationTieDecay2018].
Combinations of mechanisms can yield patterns worth examining on their own, like hierarchical structures. Hierarchies can appear at the individual or group level, and there is no clear consensus on why and how these hierarchies are formed. At the most basic level, asymmetric links may be tentative or aspirational links in a relationship in the process of formation and have not yet had time to be consolidated. However, if they are not reciprocated or dissolved, they can lead to the formation of hierarchies. Some authors suggest hierarchies could appear as a combination of reciprocity, transitivity and differential popularity [davisStructurePositiveInterpersonal1967, hollandMethodDetectingStructure1970, hollandTransitivityStructuralModels1971, hallinanCommentHollandLeinhardt1972, hallinanStructuralModelSentiment1974]. This pseudo-mechanism provides the theoretical basis behind status theory [leskovecSignedNetworksSocial2010] [leskovecPredictingPositiveNegative2010] [leskovecPredictingPositiveNegative2010]. However, there are other possible explanations behind the formation of asymmetric links. Thresholds in the definition of positive or negative versus neutral relationships across nodes may cause this asymmetry. It may also be a combination of differential popularity and a resource constraint that prevents a reciprocation of links, even if there is a willingness to do so. Finally, some findings suggest it may be a combined effect of the tendency to reciprocate and a long self-perceived centrality on the part of the nodes [anOriginsAsymmetricTies2015]. This discussion highlights, among other things, the importance of developing a comprehensive domain-specific knowledge of the system under study, which can be integrated into the mathematical framework to study the dynamics rigorously.
@@
As noted by [riveraDynamicsDyadsSocial2010], mechanisms cannot be understood in isolation, even though understanding how different mechanisms interact is as important as understanding the mechanisms themselves. They may produce overlapping or even opposing effects, making them hard to disentangle. To what extent does homophily exist in a network because actors bring themselves to contexts in which certain characteristics are shared? Is homophily still valid when I connect with my friends’ friends? How does balance operate if nodes are averse to conflict? Some interactions between mechanisms have been studied in unsigned networks. For instance, in [blockReciprocityTransitivityMysterious2015], the authors explore how transitivity and reciprocity operate together, finding a somewhat counter-intuitive result: asymmetric links are more stable when they are structurally embedded. [schaeferFundamentalPrinciplesNetwork2010] studies the joint effect of reciprocity, differential popularity and transitivity, giving rise to the concept of structural cascading.
@@
For signed networks, some studies model interactions between mechanisms in purely theoretical terms. The most notable research in this direction concerns the interaction between balance and homophily [dengInfluenceStructuralBalance2016, gorskiHomophilyBasedFew2020, phamEffectSocialBalance2020, phamBalanceFragmentationSocieties2021, phamEmpiricalSocialTriad2022]. These papers extend existing models of balance by introducing homophilic preferences in link formation.
@@
Many other papers study the competition between mechanisms at the empirical level. An early example of this is [leskovecSignedNetworksSocial2010], which compares the predictive power of balance theory and status theory. It finds that both theories make some contradicting predictions, and balance better explains triads with reciprocal links, while status is more relevant for asymmetric, hierarchical ones. This result suggests that both mechanisms operate simultaneously but affect different aspects of the structure. Other early examples are  [facchettiComputingGlobalStructural2011] and [doreianTestingTwoTheories2014], which study the influence of differential popularity on balance. The authors’ study examines the effect that the presence of positive and negative hubs has on balance. They find that when there is a negative link hub, there is a good chance that this node is involved in many balanced triangles. Two nodes connected by a positive link will have a better chance of both being connected to the hub, regardless of how balance is operating in the network. The opposite applies if the hub is positive. Finally, one of the most systematic empirical comparisons of mechanism interplay is given in [yapWhyDoesEverybody2015]. Specifically, that paper analyses a large corpus of networks and quantifies the percentage of the data’s variance explained by balance theory, by homophily, and by status theory. This paper shows how it is necessary to include the combined effect of all these mechanisms to explain the observed structure.
@@
While most of the above studies take a heuristic or model-driven approach, more systematic methodologies are emerging. A promising direction is the use of Stochastic Actor-Oriented Models (SAOMs) [snijdersStatisticalModelsSocial2011], which enable statistical inference about multiple interacting mechanisms from longitudinal data. Applications of SAOMs to signed networks, such as [lernerStructuralBalanceSigned2016, huitsingVictimsBulliesTheir2014, rambaranDevelopmentAdolescentsFriendships2015, bergerCompetitionEnvySnobbism2013, toroslu2022avoidance, kros2021avoidance], offer a rigorous framework for testing how mechanisms operate jointly.
@@ 5.3 Co-evolving dynamics
The first coevolving model of opinion dynamics with signed links was introduced in [saeedianAbsorbingPhaseTransition2019]. In this work, each node holds a binary opinion, and each link can be positive or negative. A connection is satisfying when a positive link joins nodes with the same opinion or a negative link joins nodes with opposite opinions; otherwise, either one node flips its opinion or the link changes sign to restore consistency. Through this mutual adjustment, the system evolves toward either global consensus (with only positive links) or polarisation into two antagonistic groups connected by negative edges. In a follow-up paper [saeedianAbsorbingstateTransitionCoevolution2020], the authors allowed unsatisfying links to rewire instead of merely changing sign, so that disagreeing nodes could remove hostile links and form new friendly ones. When this rewiring mechanism dominates, the network fragments into two cohesive factions; otherwise, the network is driven toward full consensus.
A significant fraction of coevolving models is based on Hamiltonian-based approaches. In particular, a series of works [phamEffectSocialBalance2020, phamBalanceFragmentationSocieties2021, phamEmpiricalSocialTriad2022, galesicExperimentalEvidenceConfirms2025] studies Hamiltonians of the form:
@@
where $A$ is the signed adjacency matrix, and $s_{i}\in\{-1,+1\}^{d}$ are vectors containing the agents’ binary opinions on multiple topics [^28] Hamiltonian formulations have also proved useful for the study of the dynamics of signed networks, without node-level features (see Section 5.2.1), and for opinion dynamics on static signed networks [liBinaryOpinionDynamics2019, Baldassarri2023].. This equation encodes two mechanisms driving the evolution of stress in the system: the correspondence between connections and opinion similarity and structural balance. Agents seek to reduce their social stress by maximising these two quantities, i.e. by minimising $H$. In the model of [phamEffectSocialBalance2020], the edges evolve alongside opinions to reduce social stress, while in [phamBalanceFragmentationSocieties2021, phamEmpiricalSocialTriad2022], they are determined directly by the similarity of agents’ opinions. In both cases, the system exhibits a first-order transition between polarised and unpolarised states depending on the model parameters and the network topology. Perhaps the most important result is that, even when removing the explicit need for structural balance, i.e., the second term in Eq. (12), dyadic interactions modelled by the first term of $H$ are sufficient to entail structural balance [phamEmpiricalSocialTriad2022]. These theoretical results are validated on a real-life network extracted from an online multiplayer game, and through a participant experiment [galesicExperimentalEvidenceConfirms2025]. In particular, the dyadic interaction model matches better with empirical observations.
@@
Beyond Ising-like models, another central contribution is by Schweighofer et al. [schweighoferWeightedBalanceModel2020, schweighoferRaisingSpectrumPolarization2024], who proposed an empirically validated model to explain opinion polarisation and issue alignment in multi-dimensional opinion spaces. The system consists of agents who evaluate their affect toward others on the basis of the similarity of their multi-dimensional opinions. In other words, the nature of the signed link between any two agents is a function of their individual (signed) links with a third party, namely the political issue of interest (e.g. support or opposition to environmental policies). The authors validate this micro-level hypothesis with data from the 2016 American National Election Study (ANES). The dynamics of the model proceed as the agents repeatedly re-evaluate their relationships, then nudge their opinions closer to those of their friends and further away from those of their foes. After a transient state of consensus, the model generates polarised structures with aligned issues. These dynamics can reproduce trends of polarisation among ANES respondents over 40 years of data [schweighoferRaisingSpectrumPolarization2024].
@@
Coevolution has also been studied in the context of contagion, where the system updates both node states and link signs to reduce social tension at the same time that the contagion advances. One way to encode this coupling is by penalising both unbalanced triads (social tension) and friendly exposure between infected and non–infected nodes (spreading pressure). This coevolution typically ends either in a balanced network or in jammed local minima [saeedianEpidemicSpreadingEvolving2017]. In both cases, unfriendly links act as an immunisation that prevents infection even if some friendly cross-links remain. A natural extension adds an “alert” compartment to capture behavioural response [kalimzhanovCoEvolution_of_Viral_Processes_and_Structural_Stability_in_Signed_Social_Networks2023]. Under this assumption, the network self-organises into two clusters (one alerted and one infected) interconnected mostly by unfriendly links, shrinking the magnitude of the epidemic.

### v2

@@ 5.2.2 Data-driven approaches
A third approach to gathering empirical signed networks involves extracting data from online platforms. This strategy has gained popularity thanks to the ease of data collection and the ability to access systems involving thousands of agents. One of the earliest studies in this direction is [223], which analyses the Slashdot Zoo network but does not focus on temporal evolution. A more complete dynamic analysis appears in [371, 370, 372], which examines the multilayer network of Pardus, a massive multiplayer online game with over 300,000 nodes. The structural analysis, though static, already points to key mechanisms in the network. They determine three different ways in which positive and negative links differ: first, positive links are highly reciprocal, whereas negative links are not; second, negative links follow a power-law degree distribution, unlike positive ones; and finally, positive layers exhibit higher clustering. In addition, the hubs in the positive and negative subgraphs are uncorrelated. The network also appears more balanced than expected by chance. In their analysis of network dynamics, the authors examine the network evolution over three years. Once again, they find qualitatively different behaviour in the link dynamics depending on their sign: negative links follow preferential attachment, whereas positive ones do not. In addition, they distinguish between two types of negative links: private enemies, which are typically local and uncorrelated with balance, and public enemies, who act as negative hubs and are often jointly targeted by friends, thus playing a central role in the formation of balanced triads. A follow-up study [334] develops a generative model to simulate Pardus’ dynamics. They show that, while the dynamics of positive links are relatively autonomous, the dynamics of negative links are tightly coupled to the positive layer. Other examples of such analyses are [236, 25, 246].
@@ 5.2.3 Mechanisms governing signed network dynamics
As we have seen, many empirical analyses agree that the network evolution is shaped by mechanisms beyond structural balance. Identifying these mechanisms has proven difficult, partly because they are often specific to the system under study and partly because they interact in complex, nontrivial ways. In addition, common simplifications such as removing link directionality can obscure important aspects of the dynamics. More fundamentally, the mechanisms we have covered so far, which we will name endogenous mechanisms, are just one part of the picture. In general, we can group the factors driving network change into three types [331]. (i) Endogenous mechanisms: some links appear or disappear by the simple presence of other links in the network, so that the current structure of the network constrains and modifies its future structure. These mechanisms include balance, reciprocity, or triadic closure. (ii) Node-based attributes that are not encoded in the network structure but affect the way the network evolves over time. Examples include personality, social status, or shared interests. (iii) Contextual and external factors, coming from the environment in which the network is embedded. These include institutional settings, cultural norms, or platform algorithms.
@@
    Homophily: the tendency to form positive links with others who share similar characteristics, such as interests, beliefs, or social status [331]. This mechanism is closely related to, and often hard to separate from, influence, which is the tendency for connected individuals to become more similar over time. Homophily can have three variations: homophobia/heterophily/heterophobia, which would be the tendency to form negative/positive/negative links with those nodes with similar/different/different characteristics, respectively. These variations can be generalised with the concept of assortativity.
@@
    Reciprocity: the tendency for directed links to become bidirectional. This results in an overrepresentation of mutual links compared to a null model. Moreover, when asymmetrical links do not become reciprocal, they tend to be unstable and eventually disappear [331].
@@
    Transitivity and clustering: the tendency to complete the connection $p\rightarrow q$ given the existing connections $p\to o$ and $o\to q$. In undirected networks, this mechanism is known as clustering or triadic closure, the tendency to close open triangles. This effect seems to be cumulative and grows with the number of common neighbours shared by two nodes. [331].
@@
    Differential popularity: the tendency of networks to display heterogeneous degree distributions for both positive and negative links [331]. This mechanism includes two distinct dimensions: activity, which refers to variation in the number of outgoing links, and popularity, which refers to variation in the number of incoming links. At a dynamic level, differential popularity can appear by means of preferential attachment or detachment, a mechanism through which the rates of gain or loss of links depend on the degree of the node [31]. It should be noted that the different ideas encompassed within the generic term “differential popularity” may be related to each other, giving rise to theories such as visibility theory, popularity theory, activity theory, or karma theory. All of these possibilities are collected in [406].
@@
    Resource allocation: forming and maintaining links requires time, attention, and cognitive effort, which are all limited. This constraint shapes how many links a person can have and how these social bonds are structured [317, 375, 376]. These resource constraints shape the degree distributions for both positive and negative links, and can explain why negative links are often unstable (conflict is costly). It also helps explain why seemingly stable structures can disappear: when links dissolve, the freed resources can be reinvested in new connections, and conversely, existing links may need to be dropped to make room for others.
@@
    Memory: while many models assume Markovian dynamics (i.e., memoryless evolution), empirical evidence suggests that past interactions can influence future ones. For instance, nodes may be more likely to re-establish past positive links or avoid reconnecting after a negative link has disappeared. A similar reasoning has been suggested to influence the process of link removal [214].
Combinations of mechanisms can yield patterns worth examining on their own, like hierarchical structures. Hierarchies can appear at the individual or group level, and there is no clear consensus on why and how these hierarchies are formed. At the most basic level, asymmetric links may be tentative or aspirational links in a relationship in the process of formation and have not yet had time to be consolidated. However, if they are not reciprocated or dissolved, they can lead to the formation of hierarchies. Some authors suggest hierarchies could appear as a combination of reciprocity, transitivity, and differential popularity [78, 181, 182, 165, 162]. This pseudo-mechanism provides the theoretical basis behind status theory [232] [231] [231]. However, there are other possible explanations behind the formation of asymmetric links. Thresholds in the definition of positive or negative versus neutral relationships across nodes may cause this asymmetry. It may also be a combination of differential popularity and a resource constraint that prevents a reciprocation of links, even if there is a willingness to do so. Finally, some findings suggest it may be a combined effect of the tendency to reciprocate and a long self-perceived centrality on the part of the nodes [12]. This discussion highlights, among other things, the importance of developing a comprehensive domain-specific knowledge of the system under study, which can be integrated into the mathematical framework to study the dynamics rigorously.
@@
As noted by [331], mechanisms cannot be understood in isolation, even though understanding how different mechanisms interact is as important as understanding the mechanisms themselves. They may produce overlapping or even opposing effects, making them hard to disentangle. To what extent does homophily exist in a network because actors bring themselves to contexts in which certain characteristics are shared? Is homophily still valid when I connect with my friends’ friends? How does balance operate if nodes are averse to conflict? Some interactions between mechanisms have been studied in unsigned networks. For instance, in [39], the authors explore how transitivity and reciprocity operate together, finding a somewhat counter-intuitive result: asymmetric links are more stable when they are structurally embedded. [342] studies the joint effect of reciprocity, differential popularity, and transitivity, giving rise to the concept of structural cascading.
@@
For signed networks, some studies model interactions between mechanisms in purely theoretical terms. The most notable research in this direction concerns the interaction between balance and homophily [82, 153, 274, 312, 313]. These papers extend existing models of balance by introducing homophilic preferences in link formation.
@@
Many other papers study the competition between mechanisms at the empirical level. An early example of this is [232], which compares the predictive power of balance theory and status theory. It finds that both theories make some contradicting predictions, and balance better explains triads with reciprocal links, while status is more relevant for asymmetric, hierarchical ones. This result suggests that both mechanisms operate simultaneously but affect different aspects of the structure. Other early examples are  [113] and [98], which study the influence of differential popularity on balance. The authors’ study examines the effect that the presence of positive and negative hubs has on balance. They find that when there is a negative link hub, there is a good chance that this node is involved in many balanced triangles. Two nodes connected by a positive link will have a better chance of both being connected to the hub, regardless of how balance is operating in the network. The opposite applies if the hub is positive. Finally, one of the most systematic empirical comparisons of mechanism interplay is given in [406]. Specifically, that paper analyses a large corpus of networks and quantifies the percentage of the data’s variance explained by balance theory, by homophily, and by status theory. This paper shows how it is necessary to include the combined effect of all these mechanisms to explain the observed structure. Another recent interesting contribution based on the usage of ERGMs is [135].
@@
While most of the above studies take a heuristic or model-driven approach, more systematic methodologies are emerging. A promising direction is the use of Stochastic Actor-Oriented Models (SAOMs) [358], which enable statistical inference about multiple interacting mechanisms from longitudinal data. Applications of SAOMs to signed networks, such as [230, 194, 323, 37, 385, 219], offer a rigorous framework for testing how mechanisms operate jointly.
@@ 5.3 Co-evolving dynamics
The first coevolving model of opinion dynamics with signed links was introduced in [337]. In this work, each node holds a binary opinion, and each link can be positive or negative. A connection is satisfying when a positive link joins nodes with the same opinion or a negative link joins nodes with opposite opinions; otherwise, either one node flips its opinion or the link changes sign to restore consistency. Through this mutual adjustment, the system evolves toward either global consensus (with only positive links) or polarisation into two antagonistic groups connected by negative edges. In a follow-up paper [336], the authors allowed unsatisfying links to rewire instead of merely changing sign, so that disagreeing nodes could remove hostile links and form new friendly ones. When this rewiring mechanism dominates, the network fragments into two cohesive factions; otherwise, the network is driven toward full consensus.
A significant fraction of coevolving models is based on Hamiltonian-based approaches. In particular, a series of works [274, 312, 313, 136, 59] studies Hamiltonians of the form:
@@
where $A$ is the signed adjacency matrix, and $s_{i}\in\{-1,+1\}^{d}$ are vectors containing the agents’ binary opinions on multiple topics [^28] Hamiltonian formulations have also proved useful for the study of the dynamics of signed networks, without node-level features (see Section 5.2.1), and for opinion dynamics on static signed networks [237, 30].. This equation encodes two mechanisms driving the evolution of stress in the system: the correspondence between connections and opinion similarity and structural balance. Agents seek to reduce their social stress by maximising these two quantities, i.e. by minimising $H$. In the model of [274], the edges evolve alongside opinions to reduce social stress, while in [312, 313, 59], they are determined directly by the similarity of agents’ opinions. In both cases, the system exhibits a first-order transition between polarised and unpolarised states depending on the model parameters and the network topology. Perhaps the most important result is that, even when removing the explicit need for structural balance, i.e., the second term in Eq. (12), dyadic interactions modelled by the first term of $H$ are sufficient to entail structural balance [313]. These theoretical results are validated on a real-life network extracted from an online multiplayer game, and through a participant experiment [136]. In particular, the dyadic interaction model matches better with empirical observations.
@@
Beyond Ising-like models, another central contribution is by Schweighofer et al. [346, 345], who proposed an empirically validated model to explain opinion polarisation and issue alignment in multi-dimensional opinion spaces. The system consists of agents who evaluate their affect toward others on the basis of the similarity of their multi-dimensional opinions. In other words, the nature of the signed link between any two agents is a function of their individual (signed) links with a third party, namely the political issue of interest (e.g. support or opposition to environmental policies). The authors validate this micro-level hypothesis with data from the 2016 American National Election Study (ANES). The dynamics of the model proceed as the agents repeatedly re-evaluate their relationships, then nudge their opinions closer to those of their friends and further away from those of their foes. After a transient state of consensus, the model generates polarised structures with aligned issues. These dynamics can reproduce trends of polarisation among ANES respondents over 40 years of data [345].
@@
Coevolution has also been studied in the context of contagion, where the system updates both node states and link signs to reduce social tension at the same time that the contagion advances. One way to encode this coupling is by penalising both unbalanced triads (social tension) and friendly exposure between infected and non–infected nodes (spreading pressure). This coevolution typically ends either in a balanced network or in jammed local minima [335]. In both cases, unfriendly links act as an immunisation that prevents infection even if some friendly cross-links remain. A natural extension adds an “alert” compartment to capture behavioural response [204]. Under this assumption, the network self-organises into two clusters (one alerted and one infected) interconnected mostly by unfriendly links, shrinking the magnitude of the epidemic.